if secA + tanA = p then find the value of secA – tanA
 

Grade:10

3 Answers

Muhammad Ali
6 years ago
secA+tanA=p
adding -2tanA on both sides we get
secA+tanA-2tanA=p-2tanA
sacA-tanA=p-2tanA
for more simplification put value of tanA.
Lalit kumar
6 years ago
we know that,1+tan²A=sec²A=>sec²A-tan²A=1=(secA+tanA)(secA-tanA)=1given is secA+tanA=pthen from above equation...p(secA-tanA)=1secA-tanA=1/p ans.
Prateek Chourasia
5 years ago
We  know that sec^2A-tan^2A=1 => (secA-tanA)(secA+tanA)=1 => (secA-tanA)p=1 => secA-tanA=1/p Here is the answer . The better side of this method is that you can also find secA and tanA in terms of p seprately